8. Given log(70) = 1.8451 , find a approximation for each logarithm
a) log(7)
b) log(700)
c) log(.07)
d) log(.7)
e) log(700000)
f) log(.007)

Can you please help me out? Thanks so much in advance:)

Answer by lwsshak3(11628)

8. Given log(70) = 1.8451 , find a approximation for each logarithm
a) log(7)=log(70)-log(10)=1.8451-1=0.8451
b) log(700)=log(70)+log(10)=1.8451+1=2.8451
c) log(.07)=log(70)-log(1000)=1.8451-3=-1.1549
d) log(.7)=log(70)-log(100)=1.8451-2=-0.1549
e) log(700000)=log(70)+log(10000)=1.8451+4=5.8451
f) log(.007) =log(70)-log(10000)=1.8451-4=-2.1549